You are provided with the following information relating to V ltd

Equity and liabilities 

12% debentures (shs1000 at par) 16,000

10% preferences shares   6,250

Ordinary shares (Shs 10 par)  12,500

Retained earnings  28,125

Additional information 

i. The debentures are currently selling at Shs 950 in the market 

ii. Company paid a dividend of Shs 5.00 per ordinary share and they are expected to grow at a rate of 10% per annum.

iv. The corporation tax is 40%

Required 

Effective Cost of debt  (3 marks)

Cost of equity (3 marks)

Weighted Average cost of capital (4 marks)



1
Expert's answer
2022-03-08T01:38:02-0500

Answer

i. Effective cost of debt = debentures - current selling price

Effective cost of debt = "1000 - 950 = Sh. 50."

ii. CoE = (Next Year's Dividends per Share/ Current Market Value of Stocks) + Growth Rate of Dividends

"(1000\/950)+5 = 6.06"

iii. average weight cost of capital is calculated by multiplying the cost of each capital source (debt and equity) by its relevant weight by market value, and then adding the products together to determine the total.

"(5*950) + 1000 = 5750."